OpenStudy (anonymous):

Is it possible to obtain resonating frequencies for macro particles like chalk powder?

OpenStudy (anonymous):

I've done some research about this concept based on ultrasonics.

OpenStudy (anonymous):

sure you do.

OpenStudy (anonymous):

So can I project ultrasonic waves with resonating frequencies that vibrates dust particles?

OpenStudy (anonymous):

sure can.

OpenStudy (anonymous):

so can I go with the equation \[\frac{ 1 }{2\Pi \sqrt{LC}} =\frac{P}{2L}\sqrt{\frac{E}{ ? } \rho}\]

OpenStudy (anonymous):

I cannot say about the equation, unless you tell me how you got it. But theoretically and practically, it is possible to vibrate particles for resonance.

OpenStudy (anonymous):

looks like an electric resonance on the left

OpenStudy (anonymous):

I do not know the right hand side

OpenStudy (anonymous):

it is piezoelectric method for resonance condition

OpenStudy (anonymous):

so I thought I could create a resonating electrical frequency that can be transmitted by something

OpenStudy (anonymous):

Your idea is not clear. However, did you check the old "Crabon microphone principle"

OpenStudy (anonymous):

that's good... anything that could transmit the resonance frequency to the test area is ok for me....

OpenStudy (anonymous):

I just want to know the resonance frequency now...